OpenStudy (anonymous):

x^4-5x^2+6 divided by x^2-2 =x^2-2-2/x^-2?

OpenStudy (anonymous):

okay so what do you want to know

OpenStudy (anonymous):

is my answer correct?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

x^4-5x^2+6=(x^2-2)(x^2-3) now you can simply divide ...

OpenStudy (anonymous):

let y=x^2 to make the problem easier :)

OpenStudy (anonymous):

x^2-3?

OpenStudy (mertsj):